OpenStudy (anonymous):

R=2+2t. If you know that s is three times greater than t how could you re write the formula?

OpenStudy (calculusxy):

3t=s I think that I the formulae

OpenStudy (anonymous):

R=2+2t since s is three times greater than t so s=3t t= s/3 so R=2+2t becomes R= 2+2(s/3) i.e. R= 2+(2/3) s